OpenStudy (anonymous):

John and Maria leave town in separate cars from the same location at the same time traveling in opposite directions. John drives 55mph while maria drives 40mph. how long will it take for them to be 228 miles apart?

OpenStudy (anonymous):

h is the time. John's distance increased by 55 miles each hour. Maria's increases by 40 miles each hour. We can translate that into the equation of a line. y=55h+40h The specific value we want is 228. 228=55h+40h 228=95h h=2.4 hours

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

P (total principal) = 12000 RA (rate of stock A) = 4% RB (rate of stock B) = 5% I (total interest) = 550 IA (interest of stock A) = ??? IB (interest of stock B) = ??? These are the values

OpenStudy (anonymous):

yes i know

OpenStudy (anonymous):

i=prt IA=PA*0.04*1 IB=PB*0.05*1 IA+IB=550 0.04PA+0.05PB=550

OpenStudy (anonymous):

where PA=principal of stock A, PB=principal of stock B, etc etc

OpenStudy (anonymous):

0.04PA+0.05PB=550 PA+PB=P PA+PB=12000 PA=12000-PB and we substitute this into the original 0.04(12000-PB)+0.05PB=550 480-0.04PB+0.05PB=550 combine like terms 0.01PB=70 PB=7000 P=PA+PB 12000=PA+7000 5000=PA There you go. $5000 in the first, $7000 in the second. The magic of systems of equations.

OpenStudy (anonymous):

thanks!!!!